OpenStudy (anonymous):

Given the equation 3x + 1 = 2 + 2x – 4, use the commutative property to rearrange the terms so that like terms are next to one another. This gives the equation 3x + 1 = 2 – 4 + 2x. Then use the associative property of addition to group the like terms. This gives the equation 3x + 1 = (2 – 4) + 2x. Next, combine like terms to get the equation 3x + 1 = – 2 + 2x. Use the subtraction property of equality to subtract 2x from both sides of the equation. This gives the equation x + 1 = – 2.

OpenStudy (anonymous):

Then use the _________________________ to subtract 1 from both sides of the equation. This gives the solution x+ 1 = – 2. Then use the _________________________ to subtract 1 from both sides of the equation. This gives the solution x = –3. Therefore, given the equation 3x + 1 = 2 + 2x – 4, x is equal to –3. Which justification was left out of the paragraph proof above?

OpenStudy (anonymous):

help!!!!!!

OpenStudy (anonymous):

what?

OpenStudy (anonymous):

The first sentence is bogus. Did you mean to type it? "Then use the ________to subtract 1from both sides of the equation. This gives the solution x+1 = -2." The next blank is " the subtraction property."